# Accurate and Diverse LLM Mathematical Reasoning via Automated PRM-Guided GFlowNets

### v2

@@ 2 Background and Related Work
The core objective of GFlowNets is to learn policies that generate complete trajectories $\tau=(s_{0},s_{1},...,s_{n},s_{f})$ such that terminal states are sampled with probabilities proportional to their rewards: $P(x)\propto R(x)$. This is achieved by learning a flow function $F:\mathcal{A}\rightarrow\mathbb{R}^{+}$ satisfying two fundamental constraints:
@@
    Flow conservation: For every non-terminal, non-source state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$:
    $\sum_{(s^{\prime},s)\in\mathcal{A}}F(s^{\prime}\rightarrow s)=\sum_{(s,s^{\prime\prime})\in\mathcal{A}}F(s\rightarrow s^{\prime\prime})$ (1)
@@
    Reward matching: For every terminal state $x\in\mathcal{X}$:
    $F(x\rightarrow s_{f})=R(x)$ (2)
From this flow function, forward and backward policies can be derived as:
$P_{F}(s^{\prime}|s)=\frac{F(s\rightarrow s^{\prime})}{F(s)}\text{ and }P_{B}(s|s^{\prime})=\frac{F(s\rightarrow s^{\prime})}{F(s^{\prime})}$ (3)
where $F(s)=\sum_{(s,s^{\prime})\in\mathcal{A}}F(s\rightarrow s^{\prime})$ represents the total outflow from state $s$.
Various training objectives have been proposed to learn effective GFlowNets. The flow-matching objective [1] directly enforces the flow constraints. The detailed-balance (DB) objective [2] operates at the level of individual transitions, enforcing a local equilibrium:
$F(s)P_{F}(s^{\prime}|s)=F(s^{\prime})P_{B}(s|s^{\prime})$ (4)
The trajectory-balance (TB) objective [22] provides a more global perspective by considering complete trajectories:
$\mathcal{L}_{TB}(\tau;\theta)=\left(\log\left(\frac{Z_{\theta}\prod_{t=1}^{n+1}P^{\theta}_{F}(s_{t}\mid s_{t-1})}{R(s_{n})\prod_{t=1}^{n}P^{\theta}_{B}(s_{t-1}\mid s_{t})}\right)\right)^{2}$ (5)
where $Z_{\theta}$ is a learnable parameter representing the total reward. Additional approaches include SubTB($\lambda$) [21], which enables learning from incomplete trajectories, and weighted DB objectives [34] that account for terminal state descendants to improve training dynamics. Such objectives share similarities with hierarchical variational inference algorithms, as described in Malkin et al. [23].
GFlowNets have demonstrated remarkable success across diverse domains. Their ability to generate diverse, high-quality samples addresses a fundamental challenge in scientific discovery tasks involving astronomically large search spaces. By learning to sample diverse high-reward candidates, GFlowNets can efficiently identify promising regions of the design space while maintaining the variety needed to accommodate additional constraints not captured in the primary reward function, such as synthesis feasibility or absence of side effects. They have been successfully applied to molecule generation [14], Bayesian structure learning [7], causal discovery [24, 7], material design [5, 26], drug discovery [28], biological sequence design and editing [14, 9], scheduling [49], and combinatorial optimization problems [47]. Recent theoretical extensions have expanded GFlowNets to continuous spaces [16], stochastic environments [27], and adversarial settings [15].
Recent work has explored GFlowNets for language generation tasks [13, 36, 12], typically operating at the token level. However, token-level approaches can struggle with maintaining semantic coherence in complex reasoning tasks, where the relationship between local token choices and global solution quality can be difficult to capture. Our work addresses this limitation by adapting GFlowNets to operate at the reasoning step level, enabling more controlled and meaningful solution generation while preserving the diversity benefits that make GFlowNets attractive for creative and exploratory tasks.
3 Process Reward Model for Mathematical Reasoning Steps
We train a PRM $U(s^{\prime}|s)$ that evaluates the quality of a proposed reasoning step $s^{\prime}$ given its predecessor $s$. The PRM outputs a score between 0 and 1, representing the probability that step $s^{\prime}$ will lead to a correct solution, assuming $s$ is correct. Unlike previous work that uses binary labels [20, 19, 40], our continuous scores enable more nuanced evaluation of step quality.
3.1 Automated Data Generation via MCTS
To train the PRM without human annotations, we use MCTS coupled with a novel data augmentation strategy, based on rollout reuse and similarity grouping (details provided in subsections 3.1, 3.2, and 3.3). For each candidate step, we perform $k=96$ rollouts (similar to Luo et al. [20]) to estimate its Monte Carlo value $MC(s)$. Crucially, we only include steps in our training dataset up to the first incorrect step (where $MC(s)=0$), as subsequent steps would be built upon invalid reasoning. The steps are identified using line breaks, and the model used to generate the data is prompted to separate the steps in its reasoning using 4 few-shots prompts.
We use a binary search strategy to efficiently identify the first incorrect step within a candidate reasoning path; subsequent steps originating from this flawed logic are disregarded during evaluation. To assess the quality of a specific step $s$ identified during this search, we perform $k$ Monte Carlo rollouts ($k=96$ in our experiments to ensure representative sampling). For each rollout, the LLM generates a complete solution attempt with a temperature of 0.6 for diversity, prompted with the original question and the solution prefix ending just before step $s$. The value of the step, denoted $MC(s)$, is then calculated as the proportion of these $k$ rollouts that ultimately yield the correct final answer to the problem.
$MC(\text{s})=\frac{\text{number of successful rollouts}}{\text{total number ofrollouts}}$
All the rollouts are then placed in a pool with statistics N(s), MC(s),Q(s,r). Which are defined as follow :
  1. 1.
    N(s) : Number of visit of state s
  2. 2.
@@ 3.1 Automated Data Generation via MCTS
    Q(s,r) : state-rollout value function defined as
    $Q(\text{s,r})=\alpha^{1-MC(\text{s})}\beta^{\frac{\text{len(r)}}{L}}$
    where $\alpha,\beta\in(0,1]$ and $L>0$ are hyperparameters
This process continues until the first error in the reasoning has been detected. Once this has been done, we can then select a new rollout among all the rollouts available in the pool using their statistics and repeat the process, creating a tree. Rollouts $(s,r)$ are selected according to tree statistics : $(s,r)=argmax_{(s,r)}[Q(s,r)+U(s)]$ as defined by [20], which is based on a variant of the PUCT algorithm [32], and where:
$U(s)=c_{puct}\frac{\sqrt{\sum_{i}{N(s_{i})}}}{1+N(s)},$
where $c_{puct}$ is a constant determining the level of exploration. This strategy initially favors rollouts with low visit counts but gradually shifts preference towards those with high rollout values. Then, as we go along, we will sample rollouts that have larger and larger values in accordance with Q(s,r), This is advantageous because instead of sampling the rollouts uniformly, this strategy allows us to sample rollouts that have larger MC(s) values, which means that the stage is of fairly good quality. This strategy would make the PRM more efficient because a PRM capable of detecting errors in good quality reasoning will be more useful than a PRM capable of detecting errors in reasoning that is initially poor.
The algorithm stops if there are no more rollouts available in the pool or if a predefined limit is reached.

### v3

@@ 2 Background and Related Work
The core objective of GFlowNets is to learn policies that generate complete trajectories $\tau=(s_{0},s_{1},...,s_{n},s_{f})$ such that terminal states are sampled with probabilities proportional to their rewards: $P(x)\propto R(x)$. This is achieved through flow conservation and reward matching constraints that ensure the learned policy samples diverse, high-reward solutions rather than converging to a single optimal path.
Various training objectives have been proposed, including trajectory-balance (TB), detailed-balance (DB), and Subtrajectory Balance (SubTB) methods. Our work builds on SubTB($\lambda$) (Madan et al., 2023), which enables learning from incomplete trajectories - particularly suitable for our step-level framework.
GFlowNets have demonstrated remarkable success across diverse domains. Their ability to generate diverse, high-quality samples addresses a fundamental challenge in scientific discovery tasks involving astronomically large search spaces. By learning to sample diverse high-reward candidates, GFlowNets can efficiently identify promising regions of the design space while maintaining the variety needed to accommodate additional constraints not captured in the primary reward function, such as synthesis feasibility or absence of side effects. Recent work has explored GFlowNets for language generation tasks (Hu et al., 2023; Takase et al., 2024; Ho et al., 2024). Most notably, Takase et al. (2024) demonstrated that GFlowNet fine-tuning can generate diverse correct solutions for mathematical reasoning tasks. However, their approach operates at the token level, with states and actions defined over individual tokens: $\pi_{\theta}(Y|X)=\prod_{i}\pi_{\theta}(y_{i}|X,y_{1:i-1})$. In contrast, our work operates at the reasoning step level, where each action corresponds to generating a complete reasoning step rather than individual tokens. This step-level granularity enables more semantically meaningful control over solution generation while maintaining the diversity benefits of GFlowNets.
The step-level approach offers several advantages over token-level methods like Takase et al. (2024): (1) Semantic Coherence: Each action generates a complete, meaningful reasoning step rather than individual tokens, avoiding the semantic fragmentation that can occur in token-level generation; (2) Reward Alignment: Step-wise PRM evaluation provides more accurate reward signals than token-level approaches, where individual tokens may not reflect reasoning quality; (3) Logical Structure: Operating at the step level naturally captures the hierarchical structure of mathematical reasoning, where each step represents a logically complete inference; (4) Quality Control: The step-level granularity enables fine-grained control over reasoning quality while maintaining coherence within each step. This is particularly important for mathematical reasoning, where partial steps or token sequences may be mathematically meaningless, but complete reasoning steps represent verifiable logical progressions.
We discuss other related work in Appendix C. The relationship between our approach and Maximum Entropy RL is discussed in Appendix D.
3 Process Reward Model for Mathematical Reasoning Steps
We train a PRM $U(s^{\prime}|s)$ that evaluates the quality of a proposed reasoning step $s^{\prime}$ given its predecessor $s$. The PRM outputs a score between 0 and 1, representing the probability that step $s^{\prime}$ will lead to a correct solution, assuming $s$ is correct.
Relationship to Prior PRM Work: Process reward modeling has evolved significantly since Lightman et al. (2023) introduced step-by-step verification with human annotations. Recent automated approaches include Zhang et al. (2024), who use MCTS to infer process rewards by estimating the probability that each step leads to correct answers, and Guan et al. (2025), who employ MCTS-guided process reward models for small language models.
Our approach differs in several key aspects: (1) Data Augmentation: Unlike prior work that discards most generated rollouts, our similarity-based augmentation leverages nearly every rollout, significantly increasing data efficiency; (2) Continuous Scoring: We use continuous scores rather than binary labels (Luo et al., 2024; Lightman et al., 2023; Wang et al., 2024b), enabling more nuanced evaluation of step quality; (3) Integration: Our PRM provides reward signals compatible with various training methods such as PPO and GFlowNets. It is particularly well-suited for step-level GFlowNet training, requiring calibrated probability scores rather than just binary scores.
3.1 Automated Data Generation via MCTS
To train our PRM without human annotations, we adapt the MCTS-based data generation approach of Luo et al. (2024) with several key modifications. Following their framework, we perform Monte Carlo rollouts to evaluate step quality, but critically extend their binary scoring (correct/incorrect) to continuous values in $[0,1]$ to capture nuanced step quality gradations.
For each candidate step $s$, we estimate its Monte Carlo value $MC(s)$ by performing $k=96$ rollouts with temperature 0.6. Unlike binary evaluation schemes, our continuous scoring allows the PRM to learn fine-grained quality distinctions between reasoning steps: MC(s)=(successful rollouts from prefix ending at step s)/kMC(s)=($\text${successful rollouts from prefix ending at step }s)/k.
We use binary search to efficiently identify the first incorrect step within candidate reasoning paths. Only steps preceding the first error are included in the training dataset, as subsequent steps build upon invalid reasoning. Therefore as illustrated in Figure 3, the PRM consistently assumes that all preceding steps are correct when evaluating the current step. In other words, the PRM effectively answers the question: to what extent can this step lead towards the correct solution, under the assumption that all prior steps are valid? The key advantage of this approach is that, when the PRM is employed to provide a reward signal during PPO or GFlowNets training, it assigns rewards solely based on the quality of the step provided as input to the PRM. Consequently, earlier mistakes do not propagate to unduly penalize the model throughout the trajectory.
We applied this methodology using Qwen2.5-Math on 70,000 problems from OpenMathInstruct2 (Toshniwal et al., 2024). Complete algorithmic details and implementation specifics are provided in Algorithm 1 (Appendix E.1).
3.2 Dataset Augmentation via Rollout Reuse and Similarity Grouping
However, directly adding all rollout steps without further processing can introduce inconsistencies. As illustrated in Table 1, steps that contribute equivalently to the reasoning process can receive different $MC(s)$ values if evaluated in isolation with a limited number of rollouts ($k=1$, which would be the case if we directly evaluated each step in a rollout independently). This is because evaluating with $k=1$ can lead to noisy and unreliable step value estimations.
Table 1: Example of Similar Steps with Different Values
Steps value
Max attended 40 college courses in 2 years 0
Within 2 years, Max enrolled in 40 college courses. 1
  • If all steps within a group originated from rollouts that led to correct (resp. incorrect) final answers, all steps in the group are assigned a value of 1 (resp. 0).
  • If there is a mixture of correct and incorrect rollouts associated with the steps in a group, all steps in the group are assigned the $MC(s)$ value that was originally estimated for the MCTS step from which these rollouts were generated, as a reasonable approximation of the value for all similar steps in the group, assuming that steps similar to a step with a known $MC(s)$ are likely to have a similar probability of leading to a correct solution.
3.3 PRM Training
We fine-tuned Qwen2.5-Math (Qwen et al., 2025) as our PRM. We train the PRM to predict the probability of a step leading to a correct solution using Binary Cross-Entropy Loss (BCELoss) $L=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$, where $y_{i}$ is the true label for the $i$-th step (i.e., the $MC(s)$ value), and $\hat{y}_{i}$ is the predicted probability for the $i$-th step, given by $\hat{y}_{i}=\text{PRM}(\text{question},\text{solution\_prefix},\text{step})$.
3.4 PRM Validation
Figure 2: Accuracy of Falcon models on a subset of MATH Hard using PRM-guided search with varying numbers of proposed steps $k$. Horizontal lines indicate the baseline accuracy of unguided Falcon3-3B, 7B, and 10B models using prompt-based decoding. Solid (resp. dotted) curves represent the accuracy of Falcon3-1B, 3B, and 7B guided by our PRM (resp. Skywork-o1-Open-PRM-Qwen-2.5-7B).
We validate our PRM through three complementary analyses. First, we assess its ability to detect step-level errors by corrupting correct solution steps in systematic ways (e.g., changing numbers, or removing key reasoning components). The PRM consistently assigns lower scores to corrupted steps, demonstrating its sensitivity to reasoning errors (See Appendix E.4).
Second, we evaluate its capability to support diverse solution paths by comparing scores assigned to different valid approaches for the same problem (using the GSM8K dataset). The PRM assigns comparable scores to different valid approaches, indicating its ability to recognize multiple correct reasoning paths (See Figure 3 in Appendix E.5).
Thirdly, we implemented a guided search strategy to enhance the mathematical reasoning capabilities of LLMs using our trained PRM, similar to the approach in Snell et al. (2024). At each step of generation, the LLM proposes $k$ candidate steps (generated with a temperature of 0.8 for diversity), and the PRM selects the step with the highest score. This step is then appended to the prompt, and the process is repeated until a complete solution is generated.
Figure 2 demonstrates the effectiveness of PRM-guided search on a subset of the MATH Hard dataset. These results highlight the value of PRMs in guiding step-by-step reasoning. By selecting the most promising steps according to the PRM’s evaluation, the guided search strategy substantially enhances the accuracy of smaller models, in some cases enabling them to approach or even surpass the performance of larger unguided models on challenging mathematical reasoning tasks. Similar applications of PRMs can be found in Wang et al. (2024a; b), as well as in related “helper” models such as the Preference Process Model in Guan et al. (2025). Additional tests of the PRM are provided in Appendix E.4
4 Step-Level GFlowNet Fine-tuning for Diverse Solutions
4.1 Step-Level GFlowNet Framework
Building on Hu et al. (2023), we adapt GFlowNets to operate at the reasoning step level rather than the token level. Our key insight is that mathematical reasoning has a natural hierarchical structure where complete reasoning steps represent semantically meaningful units.
State and Action Space: A state $s$ represents a partial solution consisting of the question and all generated reasoning steps up to that point. An action corresponds to generating a complete reasoning step, transitioning from state $s$ to state $s^{\prime}$. This step-level granularity enables fine-grained control over the reasoning process while maintaining semantic coherence within each step.
Reward Structure: The reward $R(s_{n})$ for a complete reasoning trajectory is computed using our PRM scores $R(s_{n})=\prod_{i=1}^{n}U(s_{i}|s_{i-1})$, where $U(s_{i}|s_{i-1})$ is the PRM score for step $s_{i}$ given the previous partial solution $s_{i-1}$. This multiplicative structure ensures that trajectories with any low-quality steps receive low rewards, encouraging high-quality reasoning throughout. We propose in Appendix F a Bayesian interpretation of our approach.
Policy Parameterization: The forward policy $\pi(s^{\prime}|s)$ is parameterized by an LLM that generates the next reasoning step. Following Bengio et al. (2023), we incorporate a sink state $s_{f}$ to handle variable-length solutions, with $\pi(s_{f}|s)$ representing the termination probability.
4.2 Training Objective and Implementation
We adapt the Subtrajectory Balance (SubTB) loss (Madan et al., 2023) for our step-level GFlowNet framework, following Hu et al. (2023) who were the first to use GFlowNets for LLM fine-tuning and demonstrated the effectiveness of SubTB for language generation tasks.
Mathematical Formulation: The SubTB loss for our step-level framework is:
where $\lambda\in[0,1]$ is a discount factor, $R(s_{i})$ is the PRM-based reward for partial trajectory up to step $i$, and $\pi(s_{f}|s_{i})$ represents the termination probability. This formulation ensures that the learned policy satisfies the detailed balance condition across all subtrajectories, leading to proper reward-proportional sampling.
Our training procedure maintains a replay buffer $\mathcal{B}$ of size 1000 storing complete trajectories. The replay buffer prioritizes trajectories with higher rewards to accelerate learning. For each question, we (i) generate $k$ candidate reasoning paths using current policy, (ii) evaluate rewards for terminating states using PRM, (iii) Update policy using modified SubTB loss, (iv) Store successful trajectories in $\mathcal{B}$. Our approach is detailed in Algorithm 2 in Appendix G.2. We further implement a prioritized replay buffer, temperature scheduling during trajectory generation, and gradient stabilization. We detail these additions in Appendix G.3.
5 Experiments
For the main fine-tuning experiments, we used Llama3 models as the base architectures (specific sizes detailed in Appendix G.1). These models were fine-tuned using either our step-level GFlowNet approach or a PPO baseline. Both fine-tuning methods utilized reward signals from our Process Reward Model (PRM), which was pre-trained using the methodology described in Section 3 (based on the Qwen2.5-7B-math model (Qwen et al., 2025)).
PPO Baseline Implementation: Our PPO implementation follows standard practices adapted for sequential reasoning. For each training question, the LLM incrementally constructs its reasoning trajectory by generating one step at a time. At every stage of this sequential process, only the newly proposed step is subjected to evaluation: the PRM assesses this individual step and provides a step-level reward signal. This localized feedback guides the model throughout the reasoning process, ensuring that each successive step is informed by fine-grained evaluations rather than a single outcome-level supervision. The PPO objective maximizes expected cumulative rewards using the clipped surrogate objective with entropy regularization. We use identical model architectures, training data, and PRM evaluation as our GFlowNet approach to ensure fair comparison. PPO hyperparameters include: learning rate 5e-6, batch size 144, PPO clip ratio 0.2, value function coefficient 0.5, and entropy coefficient 0.01.
The GFlowNet and PPO fine-tuning was conducted on a challenging subset of 10,000 questions from the OpenMathInstruct2 dataset (Toshniwal et al., 2024) (“augmented math” category, aligning with MATH Level 5 difficulty). We evaluated the performance of the fine-tuned models on the MATH Hard benchmark (Hendrycks et al., 2021), the GSM8K benchmark (Cobbe et al., 2021), and the SAT MATH dataset (Zhong et al., 2023) to assess both in-domain accuracy and generalization capabilities.
5.1 Main Results
We evaluate our approach on three challenging mathematical reasoning benchmarks: MATH Level 5 (our training domain), GSM8K, and SAT MATH (generalization domains). Table 2 presents our main results, comparing GFlowNet fine-tuning against PPO and baseline models:
Table 2: Performance comparison on mathematical reasoning benchmarks. MATH Level 5 was used for training; GSM8K and SAT MATH evaluate generalization. Best results for each model size are bolded.
Model MATH Level 5 SAT MATH GSM8K
Llama3.2-3B-it 14.46% 65.6% 67.8%
   + PPO 15.32% 70.0% 68.4%
   + GFlowNet 17.05% 75.0% 68.5%
Llama3.1-8B-it 17.96% 81.2% 78.1%
   + PPO 18.44% 81.2% 79.1%
   + GFlowNet 18.67% 84.4% 79.0%
The empirical results validate the effectiveness of step-level GFlowNet fine-tuning across model scales and demonstrate strong generalization capabilities. The smaller Llama3.2-3B-it model shows substantial improvements with GFlowNet fine-tuning, achieving a 2.59 percentage point gain over the baseline on MATH Level 5.
Table 3: Performance comparison on mathematical reasoning benchmarks between our PRM and Skywork-o1-Open-PRM-Qwen-2.5-7B
Model SAT MATH GSM8K
Llama3.2-3B-it + GFlowNet 65.6% 67.8%
   + Our finetuned Qwen2.5-7B-math 75.0% 68.5%
   + Open-PRM-Qwen-2.5-7B 68.8% 68.5%
Generalization Analysis: The most striking results appear on out-of-domain benchmarks. On SAT MATH, GFlowNet fine-tuning achieves particularly impressive gains: 9.4 percentage points improvement for the 3B model (75.0% vs 65.6% baseline) and 3.2 percentage points for the 8B model (84.4% vs 81.2% baseline). This strong generalization can be attributed to three factors: (1) Diverse Reasoning Patterns: By training on multiple solution approaches rather than converging to a single strategy, the model develops more robust reasoning skills that transfer across problem types; (2) Step-Level Granularity: Operating at the reasoning step level captures fundamental mathematical operations that generalize across different problem formats; (3) PRM-Guided Quality: The PRM ensures that diverse solutions maintain high logical quality, preventing the degradation often seen when optimizing for diversity alone.
The consistent improvements across both model sizes indicate that our approach scales effectively, while the superior generalization compared to PPO suggests that reward-proportional sampling captures more transferable mathematical reasoning patterns than reward maximization.
Importantly, while PRM training has a computational cost, our analysis, detailed in subsection H.4, shows that GFlowNet fine-tuning itself is surprisingly efficient, requiring less training time than comparable PPO baselines.
To further demonstrate the effectiveness of our data augmentation technique, we conducted an additional fine-tuning experiment, as shown in Table 3. In this setup, we fine-tune Llama3.2-3B-it with an open-source PRM, namely Skywork-o1-Open-PRM-Qwen-2.5-7B (He et al., 2024), and compare the outcomes with those obtained using our own PRM.
5.2 Solution Diversity Analysis
To quantitatively evaluate the diversity of solutions generated by GFlowNet fine-tuning, we used a semantic similarity metric. We chose semantic embeddings over lexical measures for mathematical reasoning diversity assessment because: (1) lexical metrics like Levenshtein distance or Jaccard similarity capture only surface-level textual differences, missing semantic equivalence in mathematical expressions (e.g., “x=4x=4” vs $x=2^{2}$”), and (2) mathematical reasoning diversity requires understanding conceptual differences between solution approaches, not just word-level variations.
We measured the pairwise semantic similarity between the reasoning steps generated by each model using the “paraphrase-MiniLM-L6-v2” model (Reimers & Gurevych, 2019), a pre-trained sentence embedding model specifically designed to capture semantic relationships. Each step was encoded into a high-dimensional space using the encoding function of the model, with the output stored as tensors. The pairwise similarity between these embeddings was then computed using the cosine similarity measure, a standard metric for assessing vector similarity in semantic spaces. This metric provides a continuous score representing the semantic proximity between different reasoning steps, where lower scores indicate greater dissimilarity and, consequently, higher solution diversity. The average semantic similarity across multiple generated solutions for each model is presented in Table 4, where lower scores indicate greater diversity, and results averaged across 1,000 problems from MATH.
Table 4: Solution diversity analysis.
Model Avg. Semantic Similarity
Llama3.2-3B-it 0.80
   + PPO 0.82
   + GFlowNet 0.78
The semantic similarity scores presented in Table 4 provide compelling evidence that GFlowNet fine-tuning effectively enhances the diversity of generated mathematical solutions. This consistent reduction in semantic similarity indicates that GFlowNet-fine-tuned models are indeed generating reasoning paths that are more semantically distinct and varied compared to the other approaches. This result is consistent with the observations made during the finetuning of the LLM with GFlowNets, as shown in Figure 4(c). Specifically, we see that during training, the LLM progressively generates steps whose probabilities become increasingly aligned with the rewards provided by the PRM (i.e., the probabilities assigned by the LLM to the tokens of a step get closer to the PRM rewards for that step). Consequently, this enhanced diversity is a direct consequence of the GFlowNet training objective, which, unlike reward-maximizing RL methods like PPO, is explicitly designed to sample from a distribution proportional to the reward. A concrete example of this induced diversity is provided in Appendix I.
6 Conclusion
Limitations.
Despite these promising results, several limitations warrant discussion. The computational cost of initial PRM training through MCTS remains significant, though this is a one-time expense that enables subsequent efficient fine-tuning. The approach requires careful tuning of replay buffer strategies and similarity threshold parameters. Additionally, while our similarity-based augmentation partially addresses potential biases in automated PRM training, more sophisticated bias detection and mitigation strategies remain important areas for development.
Future Work.
Promising directions include exploring offline GFlowNet training to reduce computational costs, developing more sophisticated semantic diversity metrics tailored for mathematical reasoning, investigating the educational impact of diverse solution generation, and extending the approach to other complex reasoning domains such as program synthesis and theorem proving.
Reproducibility Statement
To ensure full reproducibility of our results, we provide comprehensive implementation details and make all necessary resources publicly available. The complete source code for our methodology, including PRM training via MCTS data generation, similarity-based data augmentation, and step-level GFlowNet fine-tuning, is available at https://github.com/Adam-yni/GFlowNets-FineTuning. Our implementation includes four main components: (1) seed dataset generation scripts for initial data collection, (2) MCTS-based training data generation as detailed in Algorithm 1, (3) similarity-based data augmentation procedures described in Section 3.2, and (4) GFlowNet fine-tuning implementation with SubTB loss as presented in Algorithm 2. All hyperparameters, including learning rates, batch sizes and Replay Buffer Size are explicitly documented in Section G.1 and provided as default values in our code. The data processing pipeline, from initial seed generation through final PRM training dataset creation, is fully automated and includes detailed logging for verification. Additionally, we provide evaluation scripts for PRM-guided search validation and comprehensive testing procedures that allow researchers to reproduce all experimental results reported in Section 5. The computational requirements, hardware specifications, and expected training times are documented in Section H.4 to facilitate resource planning for reproduction studies.
Appendix A LLM Usage
We used LLMs to aid and polish writing throughout this manuscript. Specifically, we employed large language models for grammatical refinement, clarity improvements, and ensuring consistency in technical terminology across sections. All scientific content, experimental design, results, and interpretations represent original work by the authors, with LLMs serving solely as writing assistants for language enhancement.
Appendix B Broader Impact and Applications
Our diversity-aware approach has significant implications beyond mathematical reasoning. In educational contexts, systems using our method could better evaluate student work that differs from canonical solutions but remains mathematically sound. The step-level GFlowNet principles could generalize to other sequential reasoning domains, such as coding tasks where lines of code could be treated analogously to reasoning steps, or logical proofs where intermediate deductions follow similar hierarchical structures.
Appendix C Extended Related Work
Mathematical Reasoning in LLMs.
Recent 2024 developments include multimodal approaches like MultiMath (Peng et al., 2024), and advanced training methods like Flow of Reasoning (Yu et al., 2025), which enables diverse solution generation with minimal training examples. However, critical studies like GSM-Symbolic (Mirzadeh et al., 2025) reveal fundamental limitations in current LLMs, showing performance degradation when numerical values or problem structure vary, indicating reliance on pattern matching rather than genuine reasoning.
Generative Flow Networks.
GFlowNets have been successfully applied to molecule generation (Jain et al., 2022), Bayesian structure learning (Deleu et al., 2022), causal discovery (Manta et al., 2023; Deleu et al., 2022), material design (Cipcigan et al., 2024; Nguyen et al., 2023), drug discovery (Pandey et al., 2024), biological sequence design and editing (Jain et al., 2022; Ghari et al., 2024), scheduling (Zhang et al., 2023a), and combinatorial optimization problems (Zhang et al., 2023b). Recent theoretical extensions have expanded GFlowNets to continuous spaces (Lahlou et al., 2023), stochastic environments (Pan et al., 2023), and adversarial settings (Jiralerspong et al., 2024).
Appendix D Relationship to Maximum Entropy RL
Recent theoretical work (Tiapkin et al., 2024) establishes that GFlowNets are equivalent to maximum entropy RL when the underlying DAG forms a tree structure. In sequential mathematical reasoning, this raises the question of why our approach outperforms standard PPO.
We note that the theoretical equivalence holds for maximum entropy RL (where entropy maximization is the primary objective), not standard PPO with entropy regularization. While our PPO baseline includes entropy regularization (coefficient 0.01), this serves primarily as an exploration bonus rather than implementing the full maximum entropy framework where entropy is co-optimized with reward.
The key differences are: (1) Entropy Treatment: GFlowNets naturally sample proportionally to rewards (maximum entropy principle), while our PPO uses entropy as a small regularization term; (2) Training Objective: GFlowNets directly optimize for reward-proportional sampling, while PPO maximizes expected return with entropy bonus; (3) Exploration Dynamics: The SubTB loss provides different exploration dynamics than clipped PPO objectives.
This distinction explains our empirical improvements: even when theoretically related, the implementation details and training dynamics create meaningful practical differences.
Appendix E Process Reward Model Training
E.1 MCTS Data Generation Process
Our MCTS-based data generation extends beyond the binary evaluation framework of Luo et al. (2024) to generate rich step-level training data with continuous quality scores. This section details the technical implementation and algorithmic components.
E.1.1 Step Identification and Preprocessing
Mathematical reasoning solutions are decomposed into individual steps using line breaks as delimiters. The data generation model (Qwen2.5-Math) is instructed through 4-shot prompting to structure its reasoning with clear step separations, ensuring consistent parsing across the dataset.
E.1.2 Monte Carlo Evaluation Protocol
The core evaluation mechanism assesses each step $s$ through multiple completion attempts. Starting from the solution prefix that includes all steps up to $s$, we generate $k=96$ independent rollouts using temperature sampling (T=0.6) to promote diversity. Each rollout produces a complete solution attempt, and success is determined by final answer correctness:
$MC(s)=\frac{\text{successful rollouts from prefix ending at step s}}{k}$
This continuous scoring scheme provides granular quality assessment, distinguishing between steps that consistently lead to success (MC = 1.0), those with moderate reliability (MC $\in(0,1)$), and unreliable steps (MC = 0.0).
E.1.3 Tree Search and Rollout Management
Generated rollouts are organized within an MCTS tree structure, with each node maintaining three key statistics:
@@ E.1.3 Tree Search and Rollout Management
    $N(s)$: Visit count for state $s$
@@
    $MC(s)$: Monte Carlo evaluation score
  3. 3.
@@
    $Q(s,r)$: State-rollout value function incorporating both quality and length considerations:
$Q(s,r)=\alpha^{1-MC(s)}\beta^{\frac{\text{len(r)}}{L}}$
where hyperparameters $\alpha,\beta\in(0,1]$ balance quality preference against rollout length, and $L>0$ normalizes length effects.
E.1.4 Selection Strategy and Exploration
Rollout selection follows an adapted PUCT mechanism that balances exploitation of high-quality paths with exploration of under-visited regions:
$(s,r)=\operatorname*{arg\,max}_{(s,r)}[Q(s,r)+U(s)]$
The exploration term $U(s)$ encourages diverse tree construction:
$U(s)=c_{puct}\frac{\sqrt{\sum_{i}{N(s_{i})}}}{1+N(s)}$
where $c_{puct}$ controls exploration intensity. This strategy initially favors rollouts with low visit counts but gradually shifts preference towards those with high rollout values. By prioritizing high-quality reasoning steps (higher $MC(s)$ values), we create more effective training data for the PRM compared to uniform sampling of potentially poor-quality steps.
E.1.5 Dataset Construction and Termination
The algorithm 1 constructs training examples by exploring reasoning trees until error detection through binary search. When a step receives $MC(s)=0$ (no successful rollouts), it marks the first reasoning error, and tree exploration for that path terminates. All valid steps preceding this error point are included in the final training dataset with their computed MC scores.
The complete process terminates when the rollout pool is exhausted or computational limits are reached, yielding a comprehensive dataset of step-quality pairs for PRM training.